OpenStudy (anonymous):

Determine which ordered pair is a solution of y = x2–2. (5, 21), (–5, 24), (–1, –1), (–2, 6)

OpenStudy (anonymous):

Are you sure those answer choices are right?

OpenStudy (anonymous):

yes, but there are also these (–2, 6) (5, 21) (–1, –1) (–5, 24)

OpenStudy (anonymous):

but im pretty sure they're the same...

OpenStudy (anonymous):

None of those answers are the correct choice. If the last choice were (-2,-6) it would work.

OpenStudy (anonymous):

it was (-1,-1)

OpenStudy (anonymous):

Maybe I am looking at it wrong, I was thinking PEMDAS required you to do the multiplication of X*2 first. and then subtract the 2 from that.

OpenStudy (anonymous):

You might want to check that again. I can't find any way that (-1,-1) works.

OpenStudy (anonymous):

If you did (-1*2)-2 you would get -1=0 if you did -1^2-2 you would get -1=-3 if you did -1*(2-2) you would get -1=0 so any way you look at it (right or wrong) (-1,-1) does not work.

OpenStudy (anonymous):